OpenStudy (anonymous):

Solve for the variable in the following proportion. 1/4 is to 1 1/4 as 2 is to b

OpenStudy (triciaal):

you can express 1 1/4 as 5/4 1/4 :5/4 is the same as 2:b how you change from 1/4 to 5/4 will be the same way you change from 2 to b (1/4)/(5/4) = 5 so you had to multiply 1/4 by 5 to get 5/4 so multiply (2) by 5 to get b; b=10

OpenStudy (triciaal):

alternate (1/4)/(5/4) = (2)/(b) multiply across the = (1/4)(b) = (2)(5/4) multiply each side by 4 b = 10